# Efficient Visual State Space Model for Image Deblurring

### v1

@@ Abstract
Convolutional neural networks (CNNs) and Vision Transformers (ViTs) have achieved excellent performance in image restoration. ViTs typically yield superior results in image restoration compared to CNNs due to their ability to capture long-range dependencies and input-dependent characteristics. However, the computational complexity of Transformer-based models grows quadratically with the image resolution, limiting their practical appeal in high-resolution image restoration tasks. In this paper, we propose a simple yet effective visual state space model (EVSSM) for image deblurring, leveraging the benefits of state space models (SSMs) to visual data. In contrast to existing methods that employ several fixed-direction scanning for feature extraction, which significantly increases the computational cost, we develop an efficient visual scan block that applies various geometric transformations before each SSM-based module, capturing useful non-local information and maintaining high efficiency. Extensive experimental results show that the proposed EVSSM performs favorably against state-of-the-art image deblurring methods on benchmark datasets and real-captured images.
@@ 1 Introduction
Image deblurring aims to restore sharp images from blurry ones, attracting much attention due to the popularization of various cameras and hand-held imaging devices. This task is challenging as only the blurred images are available while lacking access to the blur and latent images.
Significant progress has been made due to the development of deep convolutional neural networks (CNNs) [40, 11, 5, 53, 3]. However, as the main operation in CNNs, the convolution is spatially invariant and spatially local. This does not capture the spatially variable properties of the image contents and cannot explore non-local information that is beneficial for deblurring.
@@
Recently, state space models (SSMs) [16, 15] have demonstrated significant potential in modeling long-range dependencies for natural language processing (NLP) tasks with linear or near-linear computational complexity. The improved SSM, specifically Mamba [13], develops a selective scan mechanism (S6) and can remember relevant information and ignore irrelevant contents while achieving linear computational complexity. This motivates us to utilize Mamba to explore useful non-local information for better image deblurring efficiently. However, as Mamba is designed for handling one-dimensional (1D) sequences, it requires first flattening the image data to a 1D image sequence if straightforwardly applying Mamba to visual tasks [12, 7]. This disrupts the spatial structure of the image, making it difficult to capture local information from various adjacent pixels. Several approaches adopt a multi-direction scan mechanism to utilize the state space model in visual applications [27, 17, 37]. However, the multi-direction scan mechanism significantly increases the computational cost.
@@
In this paper, we propose an effective and efficient visual state space model for image deblurring. We find that existing visual state space models mostly adopt several fixed-direction scanning for feature extraction, which may not explore non-local information adaptively and lead to higher computational costs, We thus develop a simple yet effective scanning strategy that captures non-local spatial information while maintaining low computational costs. Specifically, we only scan the input feature in one direction but employ a simple geometric transformation before each scan, which effectively and adaptively explores useful information with a minimal increase in computational costs.
The main contributions are summarized as follows. First, we propose a simple yet effective visual state space model that efficiently restores high-quality images. Second, we develop an effective and efficient scanning strategy that captures non-local spatial information while maintaining low computational costs. Finally, we quantitatively and qualitatively evaluate the proposed method on benchmark datasets and real-world images and show that it performs effectively and efficiently against state-of-the-art methods.
@@ 2 Related Work
In recent years, significant progress has been made in image deblurring through deep CNN-based methods [31, 40, 11, 55, 53, 5, 3]. In [31], a deep CNN is proposed based on a multi-scale framework that directly estimates clear images from blurry ones. Tao et al. [40] introduce an effective scale recurrent network to enhance the utilization of information from each scale within the multi-scale framework. Furthermore, Gao et al. [11] propose a selective network parameter-sharing method to enhance further the methods proposed in [31, 40]. Additionally, generative adversarial networks (GANs) have been widely applied in image deblurring [22, 23], aiming to enhance the quality of deblurred results by generating realistic and sharp images.
Due to the limited improvement in performance with the utilization of additional scales, Zhang et al. [55] propose an effective network that adopts a multi-patch strategy for image deblurring. The deblurring process is executed step-by-step, enabling the network to refine the output progressively. To further exploit the features extracted at different stages, Zamir et al. [53] introduce a cross-stage feature fusion technique, aiming to enhance the overall performance of the deblurring method. To mitigate the problem of high computational cost associated with multi-scale frameworks, Cho et al. [5] present a multi-input and multi-output network architecture, reducing the computational burden while maintaining the deblurring performance. Chen et al. [3] analyze the baseline modules and propose simplified versions to improve the efficiency of image restoration. However, as the convolution operation is spatially invariant and spatially local, it cannot effectively model the global and spatially-variant information, limiting its ability to achieve better image restoration.
@@
As Transformers can establish long-range dependencies and effectively model global information, it has achieved significant advancements in various high-level vision tasks such as image classification [28, 47], object detection [1, 56], and semantic segmentation [57, 51]. Researchers have extended its application to image super-resolution [26], image deblurring [54, 21], and image denoising [2, 48]. The Transformer’s self-attention mechanism requires quadratic computational complexity, which is unacceptable for image restoration tasks with high-resolution images. To reduce the computational complexity of Transformers, Zamir et al. [54] propose an efficient Transformer model that computes scaled dot-product attention in the feature depth domain. Tsai et al. [42] simplify the self-attention calculation by constructing intra- and inter-strip tokens to replace global attention. Wang et al. [48] introduce a Transformer based on a UNet architecture that applies non-overlapping window-based self-attention for single image deblurring. Kong et al.  [21] propose a frequency domain-based Transformer and achieve state-of-the-art results. Although these approaches employ various strategies to reduce computational complexity, they cannot characterize long-range dependencies and non-local information effectively. In contrast, we develop an efficient visual state space model that can explore useful non-local information with low computational costs.
@@
State Space Models (SSMs) have been a time series analysis and modeling cornerstone for decades. Recent methods [16, 15] have adopted SSMs to capture long-range dependencies for sequence modeling. SSM-based methods can be computed efficiently using recurrence or convolution, with linear or near-linear computational complexity. Gu et al. [14] propose a framework (HiPPO) that poses the abstraction of optimal function approximation concerning time-varying measures. The method [15] develops a linear state-space layer to handle long-range dependency problems. To address the issue of high computational and memory requirements induced by state representation, S4 [16] proposes a method of normalizing parameters into a diagonal structure. Furthermore, Mamba [13] introduces a selective scanning layer with dynamic weights, showcasing significant potential in natural language processing. To apply SSMs for visual tasks, recent methods [27, 37, 17] adopt multi-direction scanning strategies, which will increase the computational cost. In contrast, we propose an efficient visual scan block that employs a geometric transformation before each scan to achieve non-local information exploration with high efficiency.
@@ 3.1 Overall architecture
As shown in Figure 1, the overall architecture of the proposed efficient visual state space model is based on a hierarchical encoder-decoder framework [21]. Given a blurry image $I_{blur}\in\mathbb{R}^{H\times W\times 3}$, we first employ a $3\times 3$ convolution layer to obtain the shallow feature $F_{s}\in\mathbb{R}^{H\times W\times C}$, where $H\times W$ denotes the spatial dimension, and $C$ is the number of feature channels. Then, the shallow feature $F_{s}$ is put into a 3-level symmetric encoder-decoder network. The encoder/decoder at each level is composed of several efficient vision state space (EVSS) modules (see Section 3.2). For the encoder/decoder at level-$l$, the input feature is progressively processed by each EVSS module and the intermediate feature ${F^{l}_{enc}}$ / ${F^{l}_{dec}}$ $\in\mathbb{R}^{\frac{H}{2^{l-1}}\times\frac{W}{2^{l-1}}\times{2^{l-1}}C}$ ($l=1,2,3$ in this work) is generated. We then employ the bilinear interpolation and $1\times 1$ convolution to achieve upsampling and downsampling and add the skip connection between the encoder and decoder at each level. Finally, a $3\times 3$ convolution layer is applied to the feature ${F^{3}_{dec}}$ to generate the residual image $R\in\mathbb{R}^{H\times W\times 3}$.
@@ 3.2 Efficient vision state space module
A state space model is a mathematical framework commonly used in time series analysis and control systems. The state equation describes the evolution of an underlying system over time, representing the relationship between the system’s hidden states and their temporal dynamics. The input signal $x(t)$ is mapped to the output response $y(t)$ through the hidden state $h(t)$. It is typically modeled as a set of first-order difference or differential equations:
@@
Based on (3), Mamba [13] proposes a mechanism of selective scanning (S6) to achieve input-dependent weights and linear computational complexity simultaneously. Using a state space model for NLP tasks poses no issue since natural language data is inherently a causal sequence. However, visual tasks present significant challenges because visual data is fundamentally non-sequential and contains spatial information such as local textures and global structures. As S6 is a recursive process, when processing an input at the current timestep $t$, it can only utilize information from previous timesteps and not consider information from future timesteps.
@@
A straightforward approach to solve this problem is to scan the visual data in different directions (e.g., forward and backward). However, this strategy increases the computational cost by a significant factor. For example, the computational cost of the scanning in VMamba [27] is 4×4$\times$4 × higher than that of Mamba [13] due to its strategy of performing bidirectional scanning in the longitudinal and transverse directions. Then, a natural question is whether it can reasonably process visual data using a state space model with a minimal increase in computational costs.
The answer is YES. We develop an efficient vision state space model that explores more useful information with a minimal increase in computational costs. The key is the proposed EVS block. Instead of scanning in multiple directions, we scan only in one direction and apply one geometric transformation (e.g., flip, and transposition) to the input before each scan. Due to the translation-invariant property of convolution, the geometric transformation does not affect the convolution itself but only influences the process of selective scanning. Specifically, for each EVS block, assuming that it is in the $i$-th EVSS module of the whole network, we first transpose or flip the input feature $F_{in}$ as:
@@
Here, we flip along both the horizontal and vertical axes of the feature in $Flip(\cdot)$. According to (4), the image feature will be automatically restored to the original spatial structure after every 4 EVSS modules. In particular, if the total number of EVSS modules in our network is not divisible by 4, we can restore the original spatial structure by applying the corresponding inverse transformations, as both flip and transposition are reversible. In this way, our EVSS module effectively solves the above question, avoiding any additional computational burden other than efficient geometric transformations. Then, the selective scan can be formulated as follows:
$\begin{split}&X_{1},X_{2}=\mathrm{split}(\mathrm{Conv}_{1\times 1}(G))\\ &\hat{X_{1}}=\mathrm{S6}(\mathrm{Reshape}(\sigma(\mathrm{Dconv_{3\times 3}}(X_{1}))))\\ &\hat{X_{2}}=\sigma(X_{2})\\ &F_{out}=\mathrm{Conv}_{1\times 1}(\mathrm{Reshape}(\mathcal{L}(\hat{X_{1}}))\cdot\hat{X_{2}}),\end{split}$ (5)
where $\mathrm{Conv}_{1\times 1}(\cdot)$ denotes a convolutional layer with the filter size of $1\times 1$ pixel, $\mathrm{DConv}_{3\times 3}(\cdot)$ denotes a depth-wise convolutional layer with the filter size of $3\times 3$ pixels, $\mathcal{L}(\cdot)$ denotes a normalization layer, $\mathrm{split}(\cdot)$ splits the image features in the channel dimension, $\sigma$ denotes the GeLU activation, and S6 denotes the selective scanning mechanism proposed by [13]. Figure 1(b) shows the detailed network architecture.
@@
To effectively and efficiently transform the features from the EVSS module, we develop an efficient discriminative frequency domain-based FFN. The FFN part is typically the core component of deep learning models which can help the latent clear image reconstruction [54, 21]. FFTformer [21] develops a discriminative frequency domain-based FFN that adaptively determines which frequency information should be preserved. However, this increases the computational cost when performing the frequency domain-based operations. In contrast to DFFN which applies frequency-domain based operations in the middle of the FFN network, our approach is to perform frequency-domain screening on the features at the final stage of the FFN network. Figure 1(c) shows the detailed network architecture.

### v2

@@ Abstract
Convolutional neural networks (CNNs) and Vision Transformers (ViTs) have achieved excellent performance in image restoration. While ViTs generally outperform CNNs by effectively capturing long-range dependencies and input-specific characteristics, their computational complexity increases quadratically with image resolution. This limitation hampers their practical application in high-resolution image restoration. In this paper, we propose a simple yet effective visual state space model (EVSSM) for image deblurring, leveraging the benefits of state space models (SSMs) for visual data. In contrast to existing methods that employ several fixed-direction scanning for feature extraction, which significantly increases the computational cost, we develop an efficient visual scan block that applies various geometric transformations before each SSM-based module, capturing useful non-local information and maintaining high efficiency. In addition, to more effectively capture and represent local information, we propose an efficient discriminative frequency domain-based feedforward network (EDFFN), which can effectively estimate useful frequency information for latent clear image restoration. Extensive experimental results show that the proposed EVSSM performs favorably against state-of-the-art methods on benchmark datasets and real-world images.
$\dagger$\$$\dagger$\$footnotetext: Corresponding author
@@ 1 Introduction
Image deblurring aims to restore sharp images from blurry ones, which is drawing much attention due to the popularization of various cameras and handheld imaging devices. This task is challenging, as only the blurred images are available while losing access to the blur and latent images.
Significant progress has been made due to the development of deep convolutional neural networks (CNNs) [35, 11, 4, 47, 2]. However, as the main operation in CNNs, the convolution is spatially invariant and has limited receptive fields. This does not capture the spatially variant properties of the image contents and cannot explore non-local information that is beneficial for deblurring.
@@
Recently, state space models (SSMs) [16, 15] have demonstrated significant potential in modeling long-range dependencies for natural language processing (NLP) tasks with linear or near-linear computational complexity. The improved SSM, specifically Mamba [13], develops a selective scanning mechanism (S6) and can remember relevant information and ignore irrelevant content while achieving linear computational complexity. This motivates us to utilize SSMs to efficiently explore valuable non-local information for improved image deblurring. However, since Mamba is designed to handle one-dimensional (1D) sequences, it requires first flattening of the image data to a 1D image sequence if Mamba is easily applied to visual tasks [10, 6]. This disrupts the spatial structure of the image, making it difficult to capture local information from various adjacent pixels. Several approaches adopt a multi-direction scan mechanism to utilize the state space model in visual applications [24, 17, 32]. However, the multi-direction scan mechanism significantly increases the computational cost.
@@
In this paper, we propose an effective and efficient visual state space model for image deblurring. We find that existing visual state space models mostly adopt several fixed-direction scanning for feature extraction, which may not explore non-local information adaptively and lead to higher computational costs. We thus develop a simple yet effective scanning strategy that captures non-local spatial information while maintaining low computational costs. Specifically, we only scan the input feature in one direction, but employ a simple geometric transformation before each scan, which effectively and adaptively explores useful information with a minimal increase in computational costs. In addition, we note that recent approaches based on the Mamba framework typically derive the parameters $B$, $C$, and $\Delta$ using a purely linear transformation of the input features. This results in these parameters encoding identical spatial information, as they are all derived through linear transformations of the same input features. Consequently, this limits the model’s ability to capture diverse spatial patterns, which is detrimental to the subsequent scanning process. To overcome this, we introduce a 1D depth-wise convolution operation. This operation enables the model to capture both global patterns and local details, allowing the SSMs to focus on a broader range of spatial information and thereby enhancing their ability to capture fine spatial details. Moreover, due to the presence of the geometric transformation we propose, one-dimensional convolutions can also aggregate information from the original 2D input.
Since SSMs primarily focus on global information, we propose an efficient discriminative frequency domain-based FFN (EDFFN) to enhance local details. The original DFFN [19] learns a quantization matrix in the frequency domain mid-way through the FFN to adaptively determine which frequency information should be preserved, but this results in a significant computational cost. Since the channel dimension of the features in the middle of the FFN is often several times that of the original input (e.g., the number of middle features of DFFN [19] is three times that of the input), the computational cost becomes substantial when performing a fast Fourier transform on them. Thus, we improve the original DFFN [19] by learning the quantization matrix in the frequency domain at the end of the FFN, which greatly reduces the computational cost of the network.
The main contributions are summarized as follows.
  • We propose a simple yet effective visual state space model that efficiently restores high-quality images. Compared to existing methods based on SSM, our approach can efficiently restore blurred images with only a quarter of the computational cost.
  • We propose an efficient discriminative frequency domain-based FFN that addresses the high computational cost issue of the previous frequency domain-based FFN, achieving a significant reduction in running time without compromising performance.
  • We quantitatively and qualitatively evaluate the proposed method on benchmark datasets and real-world images and show that it performs effectively and efficiently against state-of-the-art methods.
@@ 2 Related Work
In recent years, significant progress has been made in image deblurring using deep CNN-based methods [26, 35, 11, 49, 47, 4, 2]. In [26], a deep CNN is proposed based on a multi-scale framework that directly estimates clear images from blurry ones. Tao et al. [35] introduce an effective scale recurrent network to enhance the utilization of information from each scale within the multi-scale framework. Zhang et al. [49] propose an effective network that adopts a multi-patch strategy for image deblurring. The deblurring process is executed step by step, enabling the network to refine the output progressively. To further exploit the features extracted at different stages, Zamir et al. [47] introduce a cross-stage feature fusion technique, aiming to improve the overall performance of the deblurring method. To mitigate the problem of high computational cost associated with multi-scale frameworks, Cho et al. [4] present a multi-input and multi-output network architecture, reducing the computational burden while maintaining the deblurring performance. Chen et al. [2] analyze the baseline modules and propose simplified versions to improve the efficiency of image restoration. However, since the convolution operation is spatially invariant and spatially local, it cannot effectively model the global and spatially variable information, limiting its ability to achieve better image restoration.
@@
Transformers can establish long-range dependencies and effectively model global information, thus researchers have extended its application to image super-resolution [23], image deblurring [48, 19], and image denoising [1, 43]. The self-attention mechanism in Transformer requires quadratic computational complexity, which is unacceptable for image restoration tasks with high-resolution images. To reduce the computational complexity of Transformers, Zamir et al. [48] propose an efficient Transformer model that computes scaled dot-product attention in the feature depth domain. Tsai et al. [37] simplify the self-attention calculation by constructing intra- and inter-strip tokens to replace global attention. Wang et al. [43] introduce a Transformer based on a UNet architecture that applies non-overlapping window-based self-attention for single image deblurring. Kong et al.  [19] propose a frequency domain-based Transformer and achieve state-of-the-art results. Although these methods employ various strategies to reduce computational complexity, they cannot characterize long-range dependencies and non-local information effectively. In contrast, we develop an efficient visual state space model that explores useful non-local information with low computational costs.
@@
State Space Models (SSMs) have been a time series analysis and modeling cornerstone for decades. Recent methods [16, 15] have adopted SSMs to capture long-range dependencies for sequence modeling. SSM-based methods can be computed efficiently using recurrence or convolution, with linear or near-linear computational complexity. Gu et al. [14] propose a framework (HiPPO) that offers the abstraction of the optimal function approximation in terms of time-varying measures. The method [15] develops a linear state-space layer to handle long-range dependency problems. To address the issue of high computational and memory requirements induced by state representation, S4 [16] proposes to normalize parameters into a diagonal structure. Furthermore, Mamba [13] introduces a selective scanning layer with dynamic weights, showcasing significant potential in natural language processing. To apply SSMs for visual tasks, recent methods [24, 32, 17, 41] adopt multi-direction scanning strategies, which will increase the computational cost. In contrast, we propose an efficient visual scan block that employs a geometric transformation before each scan to achieve non-local information exploration with high efficiency.
@@ 3.1 Overall architecture
As shown in Figure 1, the overall architecture of the proposed EVSSM is based on a hierarchical encoder-decoder framework [19]. Given a blurry image $I_{blur}\in\mathbb{R}^{H\times W\times 3}$, we first employ a $3\times 3$ convolutional layer to obtain the shallow feature $F_{s}\in\mathbb{R}^{H\times W\times C}$, where $H\times W$ denotes the spatial dimension and $C$ is the number of feature channels. Then, the shallow feature $F_{s}$ is put into a 3-level symmetric encoder-decoder network. The encoder/decoder at each level is composed of several efficient vision state space (EVSS) modules (see Section 3.2). For the encoder/decoder at level-$l$, the input feature is progressively processed by each EVSS module and the intermediate feature ${F^{l}_{enc}}$ / ${F^{l}_{dec}}$ $\in\mathbb{R}^{\frac{H}{2^{l-1}}\times\frac{W}{2^{l-1}}\times{2^{l-1}}C}$ ($l=1,2,3$ in this work) is generated. We then employ bilinear interpolation and $1\times 1$ convolution to achieve upsampling and downsampling and add the skip connection between the encoder and decoder at each level. Finally, a $3\times 3$ convolutional layer is applied to the feature ${F^{3}_{dec}}$ to generate the residual image $R\in\mathbb{R}^{H\times W\times 3}$.
@@ 3.2 Efficient vision state space module
A state space model is a mathematical framework commonly used in time series analysis and control systems. The state equation describes the evolution of an underlying system over time, representing the relationship between the hidden states of the system and their temporal dynamics. The input signal $x(t)$ is assigned to the output response $y(t)$ through the hidden state $h(t)$. It is typically modeled as a set of first-order difference or differential equations:
@@
Based on (3), Mamba [13] proposes a selective scanning mechanism (S6) to achieve input-dependent weights and linear computational complexity simultaneously. Using SSMs for NLP tasks poses no issue since natural language data is inherently a causal sequence. However, visual tasks present significant challenges as visual data is fundamentally non-sequential and contains spatial information such as local textures and global structures. As S6 is a recursive process, when processing an input in the current timestep $t$, it can only utilize information from previous timesteps and not consider information from future timesteps.
@@
A straightforward approach to solving this problem is to scan the visual data in different directions (e.g., forward and backward). However, this strategy increases the computational cost by a significant factor. For example, the computational cost of the scanning in VMamba [24] is 4×4$\times$4 × higher than that of Mamba [13] due to its strategy of performing bidirectional scanning in the longitudinal and transverse directions. Then, a natural question is whether it can reasonably process visual data using a state space model with a minimal increase in computational costs.
The answer is YES. We develop an efficient vision state space model that explores more useful information with a minimal increase in computational costs. The key is the proposed EVS block. Instead of scanning in multiple directions, we scan only in one direction and apply one geometric transformation (e.g., flip and transposition) to the input before each scan. Due to the translation-invariant property of convolution, the geometric transformation does not affect the convolution itself, but only influences the process of selective scanning. Specifically, for each EVS block, assuming that it is in the $i$-th EVSS module of the whole network, we first transpose or flip the input feature $F_{in}$ as:
@@
Here, we flip along both the horizontal and vertical axes of the feature in $Flip(\cdot)$. According to (4), the image feature will be automatically restored to the original spatial structure after every 4 EVSS modules. In particular, if the total number of EVSS modules in our network is not divisible by 4, we can restore the original spatial structure by applying the corresponding inverse transformations, as both flip and transposition are reversible.
Recent approaches based on the Mamba framework typically derive the parameters $B$, $C$, and $\Delta$ in a purely linear fashion from the input features. These parameters share identical spatial information, differing only along the channel dimension. Inspired by previous works on transformers, which effectively capture spatial dependencies by characterizing local information before computing self-attention, we propose a simple yet effective improvement. Specifically, after the linear projection layer, we apply a 1D depth-wise convolution with a kernel size of 7 to each of $B$, $C$, and $\Delta$. This allows our model to incorporate spatial context when deriving these parameters, thereby enriching their representation and enabling them to better capture local and global dependencies within the input data. Specifically, due to the incorporation of geometric transformations, although we apply 1D convolutions on one-dimensional sequences, the model is capable of capturing not only unidirectional information but also multidirectional information from the 2D input. This enables the capture of richer contextual features across various orientations within the 1D space.
In this way, our EVS module effectively solves the above question, avoiding any additional computational burden other than efficient geometric transformations. Then, the selective scan can be formulated as follows:
$\begin{split}&X_{1},X_{2}=\mathrm{split}(\mathrm{Linear}(G))\\ &X_{1}=\sigma(\mathrm{Dconv_{3\times 3}}(X_{1}))\\ &[\Delta,B,C]=\mathrm{DConv1D}(\mathrm{Linear}(X_{1}))\\ &\hat{X_{1}}=\mathrm{S6}(X_{1},A,B,C,D,\Delta)\\ &F_{out}=\mathrm{Linear}((\mathcal{L}(\hat{X_{1}}))\cdot\sigma(X_{2})),\end{split}$ (5)
[…]
@@
SSM primarily focuses on global information, but for image restoration, local information is equally important. In order to better characterize local information to assist in image restoration, we develop an efficient discriminative frequency domain-based FFN. The FFN part is typically the core component of deep learning models, which can help the latent clear image reconstruction [48, 19]. FFTformer [19] develops a discriminative frequency domain-based FFN (DFFN) that adaptively determines which frequency information should be preserved. However, it is time-consuming to apply the Fast Fourier Transform (FFT) to the features, especially when the number of channels of the feature is large. In contrast to DFFN that applies FFT in the middle of the FFN network, our approach performs frequency-domain screening on the features at the final stage of the FFN network using a smaller learnable quantization matrix W (as shown in Figure 1(d)), thereby significantly reducing computational costs.